Car A is traveling on a freeway at 100km/h when it is passed by Car B. If both cars are 10km away from the end of the freeway, how fast would Car B have to be traveling to be able to beat Car A by 1 minute.
Answer by mananth(16949)

Car A speed = 100km/h
Car B speed =x
distance = 10 km
time = 1/60 hour
Time by B - Time by A= 1/60 hour
10/100 - 10/(x)=1/60
LCD = 600x
multiply all terms by 600x
60x-6000=10x
50x=6000
x=6000/50
x= 120 km/h speed of Car B